According to Statistics Canada (Share of Non-Alcoholic Beverage Market, 2008), 16% of
Canadians consume coffee, 12% - tea, 72% - others. Twenty five people have registered for
a workshop on potential health hazards associated with consumption of caffeine in food
and dietary supplements.
a) Find the probability that exactly five people from the list of participants prefer drinking
either tea or coffee. Calculate the probability manually and check your answer using Excel.
b) Find the probability that at least two people out of 25 prefer drinking coffee. Calculate
your answer manually and check your answer using Excel.
c) What is the average number of people who prefer tea or coffee?

Answer by stanbon(75887)

According to Statistics Canada (Share of Non-Alcoholic Beverage Market, 2008), 16% of Canadians consume coffee, 12% - tea, 72% - others.
Twenty five people have registered for a workshop on potential health hazards associated with consumption of caffeine in food and dietary supplements.
a) Find the probability that exactly five people from the list of participants prefer drinkin either tea or coffee. Calculate the probability manually and check your answer using Excel.
b) Find the probability that at least two people out of 25 prefer drinking coffee.
Note:: P(not coffee) = 0.84
-----------------------
P(2<= x <=25) = 1-P(0< x <1) = 1 - P(x = 0) - P(x = 1)
--------------------------------------------------------------
= 1 - (0.84)^25 - 12*0.16*0.84^24 = 1 - 0.0179 - 0.0292 = 0.95291
--------------------------------
Calculate your answer manually and check your answer using Excel.
c) What is the average number of people who prefer tea or coffee?
Since 72% prefer "other", 28% must prefer tea or coffee
Average = p*n = 0.28*12 = 3.36
------------
Comment:: My first answer to you was based on a complete misreading
of your problem.
-------
Cheers,
Stan H.
-----------